Simon Companys year-end balance sheets follow. At December 31 Current I Y Ago 2 Yrs Ago $ 31,128 $ 34,628 $ 37,520 89,700 64Complete this question by entering your answers in the tabs below. Required 1 Required 2 Required 2a Required 3 Compute the rRequired 1 Required 2 Required 2a Required 3 Compute the price-earnings ratio for each year. (Round your answers to 2 decimalAssuming Simons competitor has a price-earnings ratio of 6, which company has higher market expectations for future growth?Required 1 Required 2 Required 2a Required 3 Compute the dividend yield for each year. (Round your answers to 2 decimal place
Answer #1

Answer:

1.Return on Common Stockholders Equity = Net Income / Avg. Common Stock + Retained Earnings

Current Year = 34124 / ( 162580 + 134062 + 162580 + 137574 ) / 2 = 34124 / 298398 = 11.44%

1 Year ago = 34753 / ( 162580 + 137574 + 162580 + 78196 ) / 2 = 34753 / 270465 = 12.85%

2.Price Earnings Ratio = Market Price / Earning per share

Current Year = 28 / 2.10 = 13.33

1 Year ago = 26 / 2.14 = 12.15

2a.Both the companies have a PE ratio above 6 so both have higher market expectations for future growth.

3.Dividend Yield = Dividend per share / Market Price per share

Current Year = 0.20 / 28 = 0.714%

1 Year ago = 0.10 / 26 = 0.385%